OpenStudy (anonymous):

solve for y 2xy+9y = 8

OpenStudy (anonymous):

take y common

OpenStudy (anonymous):

2x+10y=8

OpenStudy (anonymous):

2xy+9y=8 Factor out the GCF of y from 2xy+9y. y(2x+9)=8 Divide each term in the equation by (2x+9). y=(8)/(2x+9)

OpenStudy (anonymous):

2xy+9y = 8................i 2x+10y=8................ii

OpenStudy (anonymous):

IS that the question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2xy+9y = 8..............i y(2x+9)=8 y=8/(2x+9).............iii

OpenStudy (anonymous):

Now substitute iii in ii

OpenStudy (anonymous):

is the final answer: y= 4/5 - 1/5x ?

OpenStudy (anonymous):

\[y=\frac{ 4 }{ 5 }-\frac{ 1 }{ 5}x\]

OpenStudy (anonymous):

Nope, u need to find the value of x and y

OpenStudy (anonymous):

2x+10y=8 2x+10*8/(2x+9)=8 2x+80/(2x+9)=8 2x(2x+9) + 80 = 8(2x+9)

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

ok thanks!